# A Computational Tsirelson’s Theorem for the Value of Compiled XOR Games

### v2

@@ 2.4 Compiled Games
A quantum strategy for a compiled game is a tuple $(\mathcal{H},|\psi\rangle,\{A_{c\alpha}\},\{B_{yb}\})$ consisting of a Hilbert space $\mathcal{H}$, an efficiently (in QPT) preparable state $|\psi\rangle\in\mathcal{H}$, operators $A_{c\alpha}=U_{c\alpha}P_{c\alpha}$ acting on $\mathcal{H}$, where $U_{c\alpha}$ is a QPT-measurable unitary and $\{P_{c\alpha}\}_{\alpha\in\Lambda}$ (where $\Lambda$ is the set of outcomes), is a QPT-measurable PVM for all $c\in\mathrm{Enc}(sk,x)$, $x\in I_{A}$, as well as QPT-measurable PVMs $\{B_{yb}\}_{b\in O_{B}}$ acting on $\mathcal{H}$ for all $y\in I_{A}$.
@@
([KLVY23, Theorem 3.2]) If $\mathcal{G}_{\mathrm{comp}}$ is a compiled nonlocal game with underlying nonlocal game $\mathcal{G}$. Then, there exists a compiled quantum strategy $S$ for $\mathcal{G}_{\mathrm{comp}}$ and a negligible function $\eta(\lambda)$ such that
@@ 2.5 Block encodings
Given binary observables that are efficiently measurable, we will also be interested in measuring their linear combinations and products in order to infer relations. To this end, we use block encodings as in [NZ23]. See [GSLW19] for a general overview of this notion. We only define it for square matrices:
@@
$tH=\left\lparen\langle 0|^{\otimes m}\otimes I\right\rparen U\left\lparen|0\rangle^{\otimes m}\otimes I\right\rparen,\qquad\text{i.e.}\qquad U=\begin{pmatrix}tH&*\\ *&*\end{pmatrix}$ )
@@
The below lemmas summarize and generalize the discussion in [NZ23] (cf. Lemmas 11–13 therein).
@@
Lemma 2.17 follows from [GSLW19, Lemma 29]. We note that it can be extended to $k=\operatorname{poly}(\lambda)$ many block encodings (if the vector $x$ can be prepared efficiently).
@@
In the remainder of this section we will discuss how to get a strong kind of non-signalling property for strategies of compiled games. We start by recalling two lemmas from [NZ23]:
@@
(Lemma 9 in [NZ23]) Let $\lambda\in\mathbb{N}$ be a security parameter. For any two efficiently (in QPT) sampleable distributions $D_{1}$, $D_{2}$ over plaintext Alice questions, for any efficiently preparable state $|\psi\rangle$, and for any QPT-measurable POVM $\{M_{\beta}\}_{\beta}$ with outcomes in $[0,1]$, there exists a negligible function $\eta^{\prime}(\lambda)$ such that
@@
(Lemma 14 in [NZ23]) Suppose we have a QPT-implementable block encoding with scale factor $\Theta(1)$ for a (not necessarily binary) observable $\mathcal{B}$, and suppose that $\|\mathcal{B}\|\leq\Theta(1)$. Then $\mathcal{B}$ is QPT-measurable up to precision $\varepsilon$ for every $\varepsilon^{-1}=\operatorname{poly}(\lambda)$, that is, there exists a QPT-measurable POVM $\{M_{\beta}\}_{\beta}$ such that for any state $\rho$ it holds
@@
The next lemma is stated more generally than [NZ23, Lemma 15–17], but the proof is essentially the same.
@@
We follow the proof of [NZ23, Lemma 15]. Suppose for the sake of contradiction that the lemma is false. Then there exists some polynomial $p(\lambda)$ such that for infinitely many $\lambda$, we have
@@
3 The macroscopic locality approach
@@
As a corollary, we will deduce a self-testing and a parallel repetition theorem for compiled XOR games. Just as we did for ordinary nonlocal games, we define the quantum bias of a compiled game as $\beta_{q}(S,\mathcal{G}_{\mathrm{comp}})=2\omega_{q}(S,\mathcal{G}_{\mathrm{comp}})-1$ and $\beta_{q}^{*}(\mathcal{G}_{\mathrm{comp}})=2\omega_{q}^{*}(\mathcal{G}_{\mathrm{comp}})-1$.
3.1 The quantum value of compiled XOR games
In order to show that the quantum value of a compiled XOR game coincides with the quantum value of the associated nonlocal game, we need the following lemma.
Lemma 3.2.
$\sqrt{\sum_{s,y\in I_{B}}G_{xs}G_{xy}V_{sy}}=\sum_{y\in I_{B}}G_{xy}\langle u_{x}|v_{y}\rangle.$ (3.1)
Proof.
Then,
since $\|\sum_{s\in I_{B}}G_{xs}v_{s}\|^{2}=\sum_{s,y\in I_{B}}G_{xs}G_{xy}V_{sy}$. Thus we have established Eq. 3.1. ∎
The next lemma is proven in the same way as [NZ23, Lemma 26].
Lemma 3.3.
Proof.
[…]
$\left(\sum_{k}p_{k}x_{k}\right)^{2}$ $\leq\sum_{k}p_{k}x_{k}^{2}$ for $\text{for }x_{k}\in\mathbb{R},p_{k}\geq 0\text{ and }\sum_{k}p_{k}=1,$ (3.3)
$(\langle\psi|O|\psi\rangle)^{2}$ $\leq\langle\psi|O^{2}|\psi\rangle$ for quantum states |$\text{for quantum states }|\psi\rangle\text{ and observables }O.$ (3.4)
Define $\mathrm{B}_{x}=\sum_{y}G_{xy}B_{y}$. From Eq. 3.2, we deduce
Note that the post-measurement states $|\psi_{c\alpha}\rangle$ are not normalized and it holds $\sum_{\alpha}\||\psi_{c\alpha}\rangle\|^{2}=1$. Therefore, using Eq. 3.3, we get
from Eq. 3.4. ∎
We are now ready to prove the main result of this section.
Proof of Theorem 3.1.
By Lemma 3.3, we get
$\beta_{q}(S,\mathcal{G}_{\mathrm{comp}})$ $\leq\sum_{x}\sqrt{\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x)}\sum_{\alpha}\langle\psi_{c\alpha}|\left(\sum_{y}G_{xy}B_{y}\right)^{2}|\psi_{c\alpha}\rangle}.$ (3.5)
Now fix any $x_{0}\in I_{A}$. For each $x\in I_{A}$, the operator $\mathcal{B}_{x}=\left(\sum_{y}G_{xy}B_{y}\right)^{2}$ has a QPT-implementable block encoding with scale factor $\Theta(1)$ by Lemmas 2.16, 2.17 and 2.18. Moreover, $\lVert\mathcal{B}_{x}\rVert=\lVert\sum_{y}G_{xy}B_{y}\rVert^{2}\leq\lparen\sum_{y}\pi(x,y)\rparen^{2}\leq 1$. Then, using Lemma 2.21 with the distributions $D_{1}=\delta_{x}$ and $D_{2}=\delta_{x_{0}}$ over Alice’s questions, we see that there exists a negligible function $\eta_{x}(\lambda)\geq 0$ such that
and hence we can upper bound Eq. 3.5 as follows:
$=\sum_{x}\sqrt{\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x_{0})}\sum_{\alpha}\langle\psi_{c\alpha}|\mathcal{B}_{x}|\psi_{c\alpha}\rangle}+\eta^{\prime}(\lambda),$ (3.6)
$\sqrt{\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x_{0})}\sum_{\alpha}\langle\psi_{c\alpha}|\mathcal{B}_{x}|\psi_{c\alpha}\rangle}=\sqrt{\sum_{s,y}G_{xs}G_{xy}V_{sy}}=\sum_{y}G_{xy}\langle u_{x}|v_{y}\rangle.$ (3.7)
This follows from Lemma 3.2, because $V\in\mathbb{R}^{I_{B}\times I_{B}}$ is positive semidefinite (it is a nonnegative linear combination of Gram matrices) and also satisfies $V_{yy}=1$. As a consequence of this and Tsirelson’s theorem (Theorem 2.10), we obtain that
$\sum_{x}\sqrt{\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x_{0})}\sum_{\alpha}\langle\psi_{c\alpha}|\mathcal{B}_{x}|\psi_{c\alpha}\rangle}=\sum_{x,y}G_{xy}\langle u_{x}|v_{y}\rangle\leq\beta_{q}^{*}(\mathcal{G}).$ (3.8)
Thus we have proved that
Together, Theorems 2.14 and 3.1 imply that the “asymptotic” quantum value of a compiled XOR game is equal to the quantum value of the underlying XOR game. That is, if $S$ is any quantum strategy for the compiled game, then
and conversely, there always exists a strategy such that
3.2 Self-testing results for compiled XOR games
see [Slo11, Section 3]. Note that the quantity $\varepsilon$ refers to the winning probability, whereas in [Slo11, Section 3], the notion of $\varepsilon$-optimality is defined with respect to the bias.
Definition 3.4.
Lemma 3.5.
$\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(z)}\sum_{\alpha}\|\left\lparen r_{z}\mathds{1}-(-1)^{\mathrm{Dec}(\alpha)}\sum_{y}G_{zy}B_{y}\right\rparen|\psi_{c\alpha}\rangle\|^{2}\leq 2\big{\lparen}\left\lparen\lvert I_{A}\rvert+\lvert I_{B}\rvert\right\rparen\,\beta_{q}^{*}(\mathcal{G})+1\big{\rparen}\varepsilon^{\prime},$
where $\varepsilon^{\prime}=2\varepsilon+\eta^{\prime}(\lambda)$.
Proof.
Since $S$ is $\varepsilon$-optimal, and using Eqs. 3.6 and 3.8 in the proof of Theorem 3.1 (with $x_{0}=z$), we see that
$\leq\beta^{*}_{q}(\mathcal{G})+\eta^{\prime}(\lambda),$ (3.10)
From this we can draw two conclusions:
First, since the second inequality in Eq. 3.10 holds for each $x\in I_{A}$ (by Lemma 3.3), we obtain that
where $\varepsilon^{\prime}:=2\varepsilon+\eta^{\prime}(\lambda)$.
From this, it follows that
$\quad\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x)}\sum_{\alpha}\|\left\lparen r_{x}-(-1)^{\mathrm{Dec}(\alpha)}\sum_{y}G_{xy}B_{y}\right\rparen|\psi_{c\alpha}\rangle\|^{2}$
$=r_{x}^{2}-2r_{x}\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x)}\sum_{\alpha}(-1)^{\mathrm{Dec}(\alpha)}\langle\psi_{c\alpha}|\left\lparen\sum_{y}G_{xy}B_{y}\right\rparen|\psi_{c\alpha}\rangle+\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x)}\sum_{\alpha}\langle\psi_{c\alpha}|\left(\sum_{y}G_{xy}B_{y}\right)^{2}|\psi_{c\alpha}\rangle$
where we used that $r_{x}\leq 1$ in the last step.
Second, from Eq. 3.10 we see that the vector strategy $\{u_{x},v_{y}\}$ is $\varepsilon^{\prime}$-optimal. Thus it follows from Eqs. 3.11 and 3.9 that
($\left\lparen r_{x}-\sqrt{\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(z)}\sum_{\alpha}\langle\psi_{c\alpha}|\left(\sum_{y}G_{xy}B_{y}\right)^{2}|\psi_{c\alpha}\rangle}\right\rparen^{2}=\left\lparen r_{x}-\sum_{y}G_{xy}\langle u_{x}|v_{y}\rangle\right\rparen^{2}\leq 2(|I_{A}|+|I_{B}|)\beta_{q}^{*}(\mathcal{G})\varepsilon^{\prime},$
for all $x\in I_{A}$. The lemma follows from this and Eq. 3.12 for $x=z$. ∎
Theorem 3.6.
[…]
is a constant that depends only on the XOR game.
Proof.
We deduce
$\mathop{\mathbb{E}}_{c\leftarrow\mathrm{Enc}(x)}\sum_{\alpha}\|\left(r_{x}^{2}\mathds{1}-\left(\sum_{y}G_{xy}B_{y}\right)^{2}\right)|\psi_{c\alpha}\rangle\|^{2}\leq 2\Big{\lparen}(|I_{A}|+|I_{B}|)\beta_{q}^{*}(\mathcal{G})+1\Big{\rparen}2\left(r_{x}+\left(\sum_{y}|G_{xy}|\right)^{2}\right)\varepsilon^{\prime}$
from Lemma 3.5. ∎
Then, Theorem 3.6 yields
3.3 Parallel repetition of XOR games
We show an upper bound on the quantum value of the compiled parallel repetition of XOR games. Throughout this section, we will treat the number of parallel repetitions $n$ as fixed (in other words, as a constant in the security parameter). However, we remark that a similar analysis can be used to establish the soundness of any polynomial number $n$ of parallel repetitions, at the cost of a sub-exponential security guarantee for the QHE. A similar tradeoff was also shown in [KLVY23].
Definition 3.7.
While the conjunction of XOR games is not an XOR game, we can make the following definition. We use the notation $[n]=\{1,\dots,n\}$.
Definition 3.8.
We first discuss the results of [CSUU08] for nonlocal games. Clearly, any quantum strategy $S$ for $\wedge_{i=1}^{n}\mathcal{G}_{i}$ determines one for $\oplus_{i\in M}\mathcal{G}_{i}$, for any subset $M\subseteq[n]$, as follows: Alice and Bob first proceed as in strategy $S$ to obtain $a,b\in\{0,1\}^{n}$. They then output $\oplus_{i=1}^{n}a_{i}$ and $\oplus_{i=1}^{n}b_{i}$, respectively. Then one has the following relation between the quantum value of the parallel repetition and the bias of the sum games:
Lemma 3.9 ([CSUU08, Lemma 8]).
It follows from this and the nontrivial result that the bias of XOR games is multiplicative with respect to taking sums [CSUU08, Theorem 1] that one has the upper bound
The converse inequality holds for any nonlocal game, since the players can always play optimal strategies for the individual games $G_{i}$ in parallel. This establishes the parallel repetition theorem for XOR games [CSUU08, Theorem 2]:
We now use a similar argument to obtain a repetition theorem in the compiled setting. As above, we first argue that any quantum strategy $S$ for the compiled parallel repetition $(\wedge_{i=1}^{n}\mathcal{G}_{i})_{\mathrm{comp}}$ determines a quantum strategy $S_{M}$ for the compile sum games $(\oplus_{i\in M}\mathcal{G}_{i})_{\mathrm{comp}}$, for any subset $M\subseteq[n]$. To this end, we modify step 2 such that if $S$ were to output $\alpha$, $S_{M}$ instead applies $\mathrm{Eval}$ to $\alpha$ to compute $\{0,1\}^{n}\ni a\mapsto\oplus_{i\in M}a_{i}$ on the encrypted data and returns the new ciphertext $\alpha^{\prime}$. [^9] Here we assume that the QHE scheme allows for more than one homomorphic evaluation, i.e., after given an evaluated ciphertext one can keep evaluating homomorphically any circuit of one’s choice. This property is satisfied by all known QHE schemes. We omit a formal definition of this property and we refer the reader to [GHV10] for details. We also modify step 4 such that if $S$ were to output $b\in\{0,1\}^{n}$, $S_{M}$ instead outputs $\oplus_{i\in M}b_{i}$. Then the following lemma can be proved similarly as [CSUU08, Lemma 8] above:
Lemma 3.10.
Theorem 3.11.
$\omega_{q}^{*}(S,(\wedge_{i=1}^{n}\mathcal{G}_{i})_{\mathrm{comp}})\leq\left\lparen\prod_{i=1}^{n}\omega_{q}^{*}(\mathcal{G}_{i})\right\rparen+\eta(\lambda).$ ) .
Conversely, there exists a quantum strategy $S$ and a negligible function $\eta(\lambda)$ such that
$\omega_{q}^{*}(S,(\wedge_{i=1}^{n}\mathcal{G}_{i})_{\mathrm{comp}})\geq\left\lparen\prod_{i=1}^{n}\omega_{q}^{*}(\mathcal{G}_{i})\right\rparen-\eta(\lambda).$ ) .
Proof.
By Lemma 3.10, we have
Using this and Eq. 3.13, we obtain that
$\omega_{q}(S,(\wedge_{i=1}^{n}\mathcal{G}_{i})_{\mathrm{comp}})\leq\left\lparen\frac{1}{2^{n}}\sum_{M\subseteq[n]}\beta^{*}_{q}(\oplus_{i\in M}\mathcal{G}_{i})\right\rparen+\eta(\lambda)=\left\lparen\prod_{i=1}^{n}\omega_{q}^{*}(\mathcal{G}_{i})\right\rparen+\eta(\lambda),$
4 The SOS approach to bounding the compiled value
In this section, we provide an alternative proof of Theorem 3.1 using the SOS approach.
@@ 4.1 SOS’s and bounding the value of nonlocal games
We briefly review the sum of squares (SOS) approach to bounding the commuting operator value $\omega^{*}_{qc}(\mathcal{G})$ for a nonlocal game $\mathcal{G}$. The starting point is an abstract characterization of a Bell scenario. In particular, we model a two-party (non-communicating) measurement scenario with $n$ measurement settings and $m$ measurement outcomes per party using the abstract scenario algebra $\mathcal{A}_{n,m}:=\mathbb{C}[\mathbb{Z}^{*n}_{m}\times\mathbb{Z}^{*n}_{m}]$. From now on we will drop the subscripts, and just write $\mathcal{A}$ for the scenario algebra. [^10] One can define a similar scenario algebra for when Alice and Bob have different sized input and or output sets. If we let $a_{x}$ and $b_{y}$ to be the order $m$ abstract observable generators of $\mathcal{A}$, we observer that there is an alternative characterization of $\mathcal{A}$ in terms of the orthogonal spectral projections $a^{x}_{a}$ and $b^{y}_{b}$, such that
$a^{x}_{a}=\frac{1}{m}\sum^{m-1}_{k=0}(\omega^{-a}a_{x})^{k},\qquad b^{y}_{b}=\frac{1}{m}\sum^{m-1}_{k=0}(\omega^{-b}b_{y})^{k}.$ (4.1)
We briefly recall some more definitions. An element $y\in\mathcal{A}$ is a (hermitian) square if $y=x^{*}x$ for some $x\in\mathcal{A}$. Moreover, we say an element of $\mathcal{A}$ is a sum-of-squares (SOS) if it can be written as a sum of (hermitian) squares in $\mathcal{A}$. A linear functional $f:\mathcal{A}\rightarrow\mathbb{C}$ is said to be positive if $f(x^{*}x)\geq 0$ for all $x\in\mathcal{A}$. In this setting, [^11] Here we are viewing the group algebra $\mathbb{C}[\mathbb{Z}^{*n}_{m}\times\mathbb{Z}^{*n}_{m}]$ as a semi-pre-$C^{*}$-algebra, with the positive cone generated by sums of squares as described in [Oza13]. The corresponding universal $C^{*}$ algebra is canonically isomorphic to $C^{*}(\mathbb{Z}^{*n}_{m})\otimes_{max}C^{*}(\mathbb{Z}^{*n}_{m})$. an expectation on $\mathcal{A}$ (also called an abstract state) is a linear functional $\operatorname*{\mathbb{E}}:\mathcal{A}\rightarrow\mathbb{C}$, that is both positive and unital (i.e $\operatorname*{\mathbb{E}}(\mathds{1})=1$). Every commuting operator strategy $\mathcal{S}=(|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$ determines an expectation $\mathbb{E}_{\mathcal{S}}:\mathcal{A}\rightarrow\mathbb{C}$. The expectation $\mathbb{E}_{\mathcal{S}}$ is determined by first applying the *-representation mapping $a_{xa}\mapsto A_{xa}$ and $b^{y}_{b}\mapsto B_{yb}$, where the inner-product of the resulting operators with the state $|\psi\rangle\in\mathcal{H}$ is then taken in the appropriate Hilbert space. Conversely, the GNS theorem for *-algebras [Sch20, Oza13] shows that every expectation $\operatorname*{\mathbb{E}}:\mathcal{A}\rightarrow\mathbb{C}$ gives rise to a commuting operator strategy $\mathcal{S}$.
@@
is an element of the scenario algebra $\mathcal{A}$. Hence, we can realize the commuting operator value $\omega^{*}_{qc}(\mathcal{G})$ as the supremum of $p_{\mathcal{G}}$ over all expectations on $\mathcal{A}$. Furthermore, if $\nu\mathds{1}-p_{\mathcal{G}}$ is expressible as an SOS then $\nu\geq\omega^{*}_{qc}(\mathcal{G})$. For all strict upper bounds $\nu>\omega^{*}_{qc}(\mathcal{G})$ one can find an SOS decomposition for $\nu-p_{\mathcal{G}}$ [Oza13]. For the exact value $\nu=\omega^{*}_{qc}(\mathcal{G})$, it remains uncertain whether an SOS decomposition always exists [MSZ23].

### v3

@@ 2.4 Compiled Games
A quantum strategy for a compiled game is a tuple $(\mathcal{H},|\psi\rangle,\{A_{c\alpha}\},\{B_{yb}\})$ consisting of a Hilbert space $\mathcal{H}$, an efficiently (in QPT) preparable state $|\psi\rangle\in\mathcal{H}$, operators $A_{c\alpha}=U_{c\alpha}P_{c\alpha}$ acting on $\mathcal{H}$, where $U_{c\alpha}$ is a QPT-measurable unitary and $\{P_{c\alpha}\}_{\alpha\in\Lambda}$ (where $\Lambda$ is the set of outcomes), is a QPT-measurable PVM for all $c\in\mathrm{Enc}(sk,x)$, $x\in I_{A}$, as well as QPT-measurable PVMs $\{B_{yb}\}_{b\in O_{B}}$ acting on $\mathcal{H}$ for all $y\in I_{B}$.
@@
([14, Theorem 3.2]) If $\mathcal{G}_{\mathrm{comp}}$ is a compiled nonlocal game with underlying nonlocal game $\mathcal{G}$. Then, there exists a compiled quantum strategy $S$ for $\mathcal{G}_{\mathrm{comp}}$ and a negligible function $\eta(\lambda)$ such that
@@ 2.5 Block encodings
Given binary observables that are efficiently measurable, we will also be interested in measuring their linear combinations and products in order to infer relations. To this end, we use block encodings as in [19]. See [10] for a general overview of this notion. We only define it for square matrices:
@@
[…]
@@
The below lemmas summarize and generalize the discussion in [19] (cf. Lemmas 11–13 therein).
@@
Lemma 2.17 follows from [10, Lemma 29]. We note that it can be extended to $k=\operatorname{poly}(\lambda)$ many block encodings (if the vector $x$ can be prepared efficiently).
@@
In the remainder of this section we will discuss how to get a strong kind of non-signalling property for strategies of compiled games. We start by recalling two lemmas from [19]:
@@
(Lemma 9 in [19]) Let $\lambda\in\mathbb{N}$ be a security parameter. For any two efficiently (in QPT) sampleable distributions $D_{1}$, $D_{2}$ over plaintext Alice questions, for any efficiently preparable state $|\psi\rangle$, and for any QPT-measurable POVM $\{M_{\beta}\}_{\beta}$ with outcomes in $[0,1]$, there exists a negligible function $\eta^{\prime}(\lambda)$ such that
@@
(Lemma 14 in [19]) Suppose we have a QPT-implementable block encoding with scale factor $\Theta(1)$ for a (not necessarily binary) observable $\mathcal{B}$, and suppose that $\|\mathcal{B}\|\leq\Theta(1)$. Then $\mathcal{B}$ is QPT-measurable up to precision $\varepsilon$ for every $\varepsilon^{-1}=\operatorname{poly}(\lambda)$, that is, there exists a QPT-measurable POVM $\{M_{\beta}\}_{\beta}$ such that for any state $\rho$ it holds
@@
The next lemma is stated more generally than [19, Lemma 15–17], but the proof is essentially the same.
@@
We follow the proof of [19, Lemma 15]. Suppose for the sake of contradiction that the lemma is false. Then there exists some polynomial $p(\lambda)$ such that for infinitely many $\lambda$, we have
@@
3 Bounding the compiled value of an XOR game
@@
As a corollary to Theorem 3.1, in Section 4 and Section 5 we derive a self-testing and a parallel repetition theorem for compiled XOR games. Just as we did for ordinary nonlocal games, we define the quantum bias of a compiled game as $\beta_{q}(S,\mathcal{G}_{\mathrm{comp}})=2\omega_{q}(S,\mathcal{G}_{\mathrm{comp}})-1$ and $\beta_{q}^{*}(\mathcal{G}_{\mathrm{comp}})=2\omega_{q}^{*}(\mathcal{G}_{\mathrm{comp}})-1$.
@@ 3.1 SOS’s and bounding the value of nonlocal games
We briefly review the sum of squares (SOS) approach to bounding the commuting operator value $\omega^{*}_{qc}(\mathcal{G})$ for a nonlocal game $\mathcal{G}$. The starting point is an abstract characterization of a Bell scenario. In particular, we model a two-party (non-communicating) measurement scenario with $n$ measurement settings and $m$ measurement outcomes per party using the abstract scenario algebra $\mathcal{A}_{n,m}:=\mathbb{C}[\mathbb{Z}^{*n}_{m}\times\mathbb{Z}^{*n}_{m}]$. From now on we will drop the subscripts, and just write $\mathcal{A}$ for the scenario algebra. [^1] One can define a similar scenario algebra for when Alice and Bob have different sized input and or output sets. If we let $a_{x}$ and $b_{y}$ to be the order $m$ abstract observable generators of $\mathcal{A}$, we observe that there is an alternative characterization of $\mathcal{A}$ in terms of the orthogonal spectral projections $a^{x}_{a}$ and $b^{y}_{b}$, such that
$a^{x}_{a}=\frac{1}{m}\sum^{m-1}_{k=0}(\zeta^{-a}a_{x})^{k},\qquad b^{y}_{b}=\frac{1}{m}\sum^{m-1}_{k=0}(\zeta^{-b}b_{y})^{k}.$ (3.1)
Where $\zeta$ is an $m$th primitive root of unity. We briefly recall some more definitions. An element $y\in\mathcal{A}$ is a (hermitian) square if $y=x^{*}x$ for some $x\in\mathcal{A}$. Moreover, we say an element of $\mathcal{A}$ is a sum-of-squares (SOS) if it can be written as a sum of (hermitian) squares in $\mathcal{A}$. A linear functional $f:\mathcal{A}\rightarrow\mathbb{C}$ is said to be positive if $f(x^{*}x)\geq 0$ for all $x\in\mathcal{A}$. In this setting, [^2] Here we are viewing the group algebra $\mathbb{C}[\mathbb{Z}^{*n}_{m}\times\mathbb{Z}^{*n}_{m}]$ as a semi-pre-$C^{*}$-algebra, with the positive cone generated by sums of squares as described in [20]. The corresponding universal $C^{*}$ algebra is canonically isomorphic to $C^{*}(\mathbb{Z}^{*n}_{m})\otimes_{max}C^{*}(\mathbb{Z}^{*n}_{m})$. an expectation on $\mathcal{A}$ (also called an abstract state) is a linear functional $\operatorname*{\mathbb{E}}:\mathcal{A}\rightarrow\mathbb{C}$, that is both positive and unital (i.e $\operatorname*{\mathbb{E}}(\mathds{1})=1$). Every commuting operator strategy $\mathcal{S}=(|\psi\rangle,\{A_{xa}\},\{B_{yb}\})$ determines an expectation $\mathbb{E}_{\mathcal{S}}:\mathcal{A}\rightarrow\mathbb{C}$. The expectation $\mathbb{E}_{\mathcal{S}}$ is determined by first applying the *-representation mapping $a_{xa}\mapsto A_{xa}$ and $b^{y}_{b}\mapsto B_{yb}$, where the inner-product of the resulting operators with the state $|\psi\rangle\in\mathcal{H}$ is then taken in the appropriate Hilbert space. Conversely, the GNS theorem for *-algebras [23, 20] shows that every expectation $\operatorname*{\mathbb{E}}:\mathcal{A}\rightarrow\mathbb{C}$ gives rise to a commuting operator strategy $\mathcal{S}$.
@@
is an element of the scenario algebra $\mathcal{A}$. Hence, we can realize the commuting operator value $\omega^{*}_{qc}(\mathcal{G})$ as the supremum of $p_{\mathcal{G}}$ over all expectations on $\mathcal{A}$. Furthermore, if $\nu\mathds{1}-p_{\mathcal{G}}$ is expressible as an SOS then $\nu\geq\omega^{*}_{qc}(\mathcal{G})$. For all strict upper bounds $\nu>\omega^{*}_{qc}(\mathcal{G})$ one can find an SOS decomposition for $\nu\mathds{1}-p_{\mathcal{G}}$ [20]. For the exact value $\nu=\omega^{*}_{qc}(\mathcal{G})$, it remains uncertain whether an SOS decomposition always exists [17].